OpenStudy (anonymous):

limit as x->pi/2 of [tanx - 1/(x-pi/2)]

OpenStudy (anonymous):

so i had this on a quiz today and i said it was undefined

OpenStudy (anonymous):

looks like the limit will not exist. either that or you could say \(-\infty\)

OpenStudy (anonymous):

yay! so i'm right?

OpenStudy (anonymous):

i would have said that as well so i bet you got it right

OpenStudy (anonymous):

it was a quiz on l'hospital's rule

OpenStudy (anonymous):

well you cannot use l'hopital here unless you change the form, because it is not \(\frac{0}{0}\)

OpenStudy (anonymous):

and i was talking to someone who was saying that he first changed tanx to sinx/cosx then cross multiplied to make them have a common denominator, then was going to try to get a form of 0/0, but he didn't finish.

OpenStudy (anonymous):

so i wasn't sure how to go about doing that or if that would work

OpenStudy (anonymous):

i think that is too much work, but maybe you could do something like that \[\frac{\sin(x)}{\cos(x)}-1=\frac{\sin(x)-\cos(x)}{\cos(x)}\]nah too much work

OpenStudy (anonymous):

k cool.

OpenStudy (anonymous):

so i guess my follow up question is, if i split the limit into lim tanx + lim (1/(x-pi/2)), then say they're both undefined because they approach infinity and negative infinity depending on the direction you're approaching pi/2 from, does that necessarily imply that the original limit is also undefined?

OpenStudy (turingtest):

I think \(\infty-\infty\) is indeterminate, so you can't stop there

OpenStudy (phi):

If the problem is \[\lim_{x \rightarrow \frac{\pi}{2}}\tan(x)+\frac{1}{x-\frac{\pi}{2}}=0\] It takes multiple steps to get to this answer. First, it is not in the form 0/0 . we could do \[\frac{\sin(x)}{\cos(x)} + \frac{1}{x-\frac{\pi}{2}}\] \[ \frac{\sin(x)(x-\frac{\pi}{2})+\cos(x) } {\cos(x)(x-\frac{\pi}{2})}\] which is now in 0/0 form. the derivative of the numerator gives us \[ \sin(x) +x \cos(x) -\frac{\pi}{2}\cos(x)-\sin(x) = x \cos(x) -\frac{\pi}{2}\cos(x)\] the derivative of the denominator is \[ -x\sin(x)+\cos(x)+\frac{\pi}{2}\sin(x) = \cos(x)+(\frac{\pi}{2}-x)\sin(x)\] we again have 0/0 However, one more set of differentiations will give us \( \frac{0}{-2}=0\)